a body dropped from height 20mt above the ground.if the gravity disaapears 1sec after it starts fallling.then total time to hit ground taking g=10ms2
 

Grade:10

1 Answers

Aziz Alam IIT Roorkee
9 years ago
The velocity attained after 1stsecond of motion.
v=at=5m/s

Height to which the body falls,
s=1/2at^2
s=0.5*10*1=5m

The time the body will take to cover the remaining distance of 15 m,
t=15/5=3s

Hence the total time to hit the ground
t=3+1=4 seconds